OpenStudy (oswinoswald):

f(-5)=|2(x)-1| +10... I've got most of it figured out, just got stuck.

OpenStudy (anonymous):

f(-5)=|2(x)-1| +10... f(-5)=|2(-5)-1| +10... f(-5)=|-10-1| +10... f(-5)=|-11| +10. f(-5)=11 +10 f(-5)=21

OpenStudy (anonymous):

REMEMBER \[\left| -x \right| = x\]

OpenStudy (oswinoswald):

I had it up to f(-5)=|-11| +10 haha... Thank you.

OpenStudy (anonymous):

you are welcome pretty :)